OpenStudy (anonymous):

how do i express the set C={x/x+6=10} in roster form

jimthompson5910 (jim_thompson5910):

what is the solution to x+6 = 10

OpenStudy (anonymous):

it has to be in roster form

jimthompson5910 (jim_thompson5910):

yes I know, but what is the solution to x+6 = 10

OpenStudy (anonymous):

not sure i forgot

jimthompson5910 (jim_thompson5910):

x is some number x+6 = 10 is like saying "some number plus 6 equals 10"

jimthompson5910 (jim_thompson5910):

what goes in place of "some number"

OpenStudy (anonymous):

would it be 4?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so x = 4

jimthompson5910 (jim_thompson5910):

that means that C = {x | x+6 = 10} is really the same as C = { 4 }

OpenStudy (anonymous):

oh ok so how wold i write that in roster form

jimthompson5910 (jim_thompson5910):

it's already in roster form

jimthompson5910 (jim_thompson5910):

C = { 4 } is in roster form

jimthompson5910 (jim_thompson5910):

there is only one element in this set because there is only one solution to x+6 = 10

OpenStudy (anonymous):

so would it be {1,2,3,4,510} or no

jimthompson5910 (jim_thompson5910):

yes something like that...BUT...the numbers: 1, 2, 3, 5, 10 do NOT satisfy x+6 =10 only 4 does

jimthompson5910 (jim_thompson5910):

that's why the answer is simply { 4 }

jimthompson5910 (jim_thompson5910):

that's what C = {x | x+6 = 10} means C = {x | x+6 = 10} means "the set of all numbers x such that x+6 = 10" only x = 4 fits this requirement, so that's why the answer is C = { 4 }

OpenStudy (anonymous):

oh i get it thank you :)

jimthompson5910 (jim_thompson5910):

that's great, I'm glad it's clicking you're welcome

OpenStudy (anonymous):

well kinda it is

jimthompson5910 (jim_thompson5910):

which part isn't making sense?

OpenStudy (anonymous):

all of it like the homework im doing

jimthompson5910 (jim_thompson5910):

oh the other problems you mean?

OpenStudy (anonymous):

yes sorry thats what i mean

jimthompson5910 (jim_thompson5910):

that's ok, do what you can and ask any questions you have by posting a new question off to the left

OpenStudy (anonymous):

ok